Problem: In a small car factory, a car type is manufactured by hand. The workers are paid with 1 USD per item manufactured. It is found that when a worker manufactures 80 items per day, 5% of them are rejected as scrap. If the worker produces 90 items/day, then 10% of them are rejected as scrap and if he produces 100 items/day 15% is rejected as scrap. The cost of materials used is 1,25 USD per item and the rejected items cannot be used again. There are some general expenses of 25 USD per day per per worker, independent of severe fluctuations in the production. Questions: (a) At which of these three production rates the workers receive the highest income? (b) At what production rate does the company have the lowest per unit cost?
Answer #1

Assumption used: it is assumed that worker is paid only for good units manufactured

Answer (a).

AT 100 units a day with 15% rejection worker earns highest income

as shown and calculated below

Wages at different level of production
If 80 Items are produced
Income to Worker =80*95%*1 =$76
If 90 Items are produced
Income to Worker =90*90%*1 =$81
If 90 Items are produced
Income to Worker =100*85%*1 =$85

Answer (b).

At 80 units with 5% scrap rate  production rate the company has the lowest per unit cost

as calculated and shown below

Units Introduce 80 90 100
Less: Scrape units =80*5% =4 =90*10%=9 =100*15% = 15
Good Units produced 76 81 85
Cost of Production At Different level
80 Units 90 Units 100 Units
Material Used =125*80 10000 =125*90 11250 =125*100 12500
Labor cost 76 81 85
General Expenses per day 25 25 25
10101 11356 12610
Good Units Produced 76 Units 81 Units 85 Units
Cost per unit =10101/76 132.91 =11356/81 140.20 =12610/85 148.35

All the best

Hit thumbs up if satisfied

Have query please mention in comment section I ll reply

Thank You